# Soft Condorcet Optimization for Ranking of General Agents

### v2

@@ 2.4 Learning-to-Rank
Another related field is that of learning-to-rank (Liu, 2011). The canonical example is that of a search engine: a user enters a keyword (query) and the problem is to retrieve the most relevant documents in a database, ranked by relevance to the keyword. There are several algorithms that learn to rank; Google’s PageRank, which powers their search engine, is one example. Our setting can be characterized by a learning-to-rank problem with a constant keyword (or no keyword) as there is no query. An important class of statistical models in this setting is random utility models (Xia, 2019). In random utility models, assessments are perceived as some ground truth assessment plus some noise.
@@
Perturbed optimizers (Berthet et al., 2020; Blondel et al., 2020b) transform non-differentiable functions (such as sorting and ranking) into smoothed versions by adding noise; these perturbed functions can then be optimized by gradient descent. This is part of a growing effort to allow end-to-end training through discrete operators, using classical stochastic smoothing and perturbation approaches (Gumbel, 1954; Hazan et al., 2016). Included are optimal transport, clustering, dynamic time-warping and other dynamic programs (Cuturi, 2013; Cuturi and Blondel, 2017; Mensch and Blondel, 2018; Vlastelica et al., 2019; Paulus et al., 2020; Sander et al., 2023; Stewart et al., 2022) applied in fields such as computer vision, audio processing, biology, and physical simulators (Cordonnier et al., 2021; Kumar et al., 2021; Carr et al., 2021; Le Lidec et al., 2021; Baid et al., 2023; Llinares-López et al., 2023) and other optimization algorithms (Dubois-Taine et al., 2022).
3 Soft Condorcet Optimization
@@
An SCO rating is a numerical value representing an agent’s level of skill compared to other agents. SCO is closely related to several prior works: Elo (Elo, 1978), probabilistic ranking (Diaconis, 1988; Marden, 1995; Alvo and Yu, 2014), and perturbed optimizers (Blondel et al., 2020b; Berthet et al., 2020). We elaborate on these relationships in Appendix C. We now present the main ideas followed by different ways to compute the SCO ratings.
3.1 SCO Ratings and the Sigmoid Loss
@@
Let $\Pi(A)$ denote the set of permutations over alternatives $A$. We will call the data we work with “votes” to emphasize that we are working with (partial) rankings over alternatives, and borrow terminology from preferences and social choice. In particular, we view the entire dataset to be a collection of votes and so will refer to it as preference profile, []delimited-[]succeeds-or-equals[$\succeq$][ ⪰ ]. Let $v\in[\succeq]$ refer to each vote in the profile, where $v$ is a permutation over subsets of $A$, with length $|v|$. For a vote $v$, denote $v[i]$ as the alternative in position $i$ such that vote $v$ is represented as: $v[0]\succ v[1]\succ\cdots\succ v[|v|-1]$.
@@
Given some a profile []delimited-[]succeeds-or-equals[$\succeq$][ ⪰ ] (i.e., set of votes), an optimal ranking minimizes the sum of the Kendall-tau distances
$R=argmin_{R\in\Pi(A)}\sum_{v\in[\succeq]}K_{d}(v,R).$
@@
$L([\succeq],A,V,\bm{\theta})=\sum_{v\in[\succeq]}\sum_{(i,j)\in I_{2}(v)}D_{v}(\theta_{v[i]},\theta_{v[j]}),$ (3)
@@
Let []delimited-[]succeeds-or-equals[$\succeq$][ ⪰ ] be the preference profile depicted in Table 1. We now show that the main loss function (Equation (3)) leads to same value as in Figure 1. The outer sum of Equation 3 enumerates the votes, which we will assume is in the same order as listed in Figure 1. The inner sum computes the Kendall-tau distance from the vote $v$ to the ranking $R$ induced by $\bm{\theta}$ (red exes in Figure 1). For the first vote $v=A\succ B\succ C$, the discrepancy function $D$ outputs 1 two times: once with pair $(i,j)=(0,2)$ and once with pair $(i,j)=(1,2)$ because the preferences between agents $(A,C)$ and agents $(B,C)$ disagree between $\bm{\theta}$ and $v$. Hence the inner sum for the first vote is 2. Similarly for the other votes: they correspond precisely to the same values as in Figure 1. Hence, the loss function (Equation (3)) is simply computing the sum of Kendall-tau distances between the ranking induced by $\bm{\theta}$ and all the votes.
@@
$\tilde{L}([\succeq],A,V,\bm{\theta})=\sum_{v\in[\succeq]}\sum_{(i,j)\in I_{2}(v)}\tilde{D}_{v}(\theta_{v[i]},\theta_{v[j]}).$ (6)
@@
3.2 Sigmoid Loss Minimization
@@
3.2.1 Gradient Descent
Input: A preference profile []delimited-[]succeeds-or-equals[$\succeq$][ ⪰ ]
Input: An initial parameter vector: $\bm{\theta}^{0}=\left(\frac{\theta_{max}-\theta_{min}}{2}\right)\mathbf{1}$
@@ 3.2.1 Gradient Descent
2       $B\leftarrow$ $K$ votes sampled unif. from []delimited-[]succeeds-or-equals[$\succeq$][ ⪰ ]
3       Define $\nabla_{\bm{\theta}}\tilde{L}(B,\bm{\theta})$ based on Eq. (7 )
4       $\bm{\theta}^{t}\leftarrow\textsc{Proj}(\bm{\theta}^{t-1}-\alpha^{t}\nabla_{\bm{\theta}}\tilde{L}(B))$
@@
The most straight-forward way is to apply gradient descent (Hazan, 2015): update ratings by following the gradient of the sigmoid loss, as shown in Algorithm 1. After applying the gradient to the ratings on line 1, the ratings may escape the bounded constraint space so we project them back. This is a straight-forward application of standard gradient descent (Hazan, 2015, Chapter 2). A common variant is stochastic gradient descent (SGD) which estimates the gradient by sampling subsets, i.e., “batches”, of the data set, computing the gradient using the sampled batch only. The standard $\ell_{2}$ projection step, Proj, projects the ratings back to the hypercube by clipping any ratings that fall outside the valid range $[\theta_{min},\theta_{max}]$.
We denote the gradient for a subset of the votes (i.e., batch $B\subseteq[\succeq]$), $\nabla\tilde{L}_{\bm{\theta}}(B)$, as equation (6) but summed only over the votes in $B$ using the continuous $\tilde{D}_{v}$ from equation (5) where
3.2.2 Sigmoidal Programming
A different way to compute SCO ratings is via sigmoidal programming (Udell and Boyd, 2014): solve for the (soft) optimum directly, i.e., find $\bm{\theta}^{*}$ that minimizes $\tilde{L}$ defined in equation 6. Note the $\tilde{L}$ can be rewritten in terms of the number of pairwise interactions between agents, quantified in the $\mathbf{N}$ matrix:
$\tilde{L}([\succeq],A,V,\bm{\theta})$ $=$ $\sum_{a,b\in A\times A}N(a,b)\sigma(\theta_{b}-\theta_{a})$ (8)
which is a sum of sigmoidal functions $\sigma$ defined as functions which are strictly convex on domain $\theta_{b}-\theta_{a}\leq z$ and then strictly concave on $\theta_{b}-\theta_{a}\geq z$ (i.e.$z=0$). With a variable per difference in pair of ratings and appropriate constraints on variables and their feasible regions, the resulting optimization problem is known as a sigmoidal program which can be solved using a branch-and-bound algorithm (Udell and Boyd, 2014). A detailed construction is presented in Appendix B.
3.3 Fenchel-Young Loss Optimization
As described in Section 2.4, another way to learn SCO ratings is via perturbed optimizers (Berthet et al., 2020), and the Fenchel-Young loss (Blondel et al., 2020a).
This loss, beyond having gradients that can be efficiently computed (requiring only to perform sorting on perturbed coefficients of $\theta$) is differentiable and convex in $\theta$ and can thus be efficiently over a convex set. Further, if the observed ranks $y_{i}$ are indeed generated from noisy sorting of a vector of ranks $\theta_{0}$, then the loss is minimized at $\theta_{0}$ (see (Berthet et al., 2020) and the appendix for details).
3.4 Theoretical Properties
@@
$\tilde{L}([\succeq],A,V,\bm{\theta})$ $=$ $\sum_{a,b\in A\times A}N(a,b)\sigma(\theta_{b}-\theta_{a})\,,$ (9)
@@
Let $c\in A$ be the Condorcet winner. The loss $\tilde{L}$ as expressed in equation 8 is expressable in terms of a constant $K$ (that does not depend on $\bm{\theta}$) and a sum of sigmoids multiplied by coefficients from $\mathbf{M}$, by symmetry of the logistic function: $\sigma(x)=1-\sigma(-x)$. The second term is decomposable into contribution from comparisons to agent $c$, which is monotonically decreasing in $\theta_{c}$, and a sum which does not depend on $\theta_{c}$. As a result, increasing $\theta_{c}$ always decreases the loss, hence the minimum must correspond to $\theta^{*}_{c}=\theta_{max}$. ∎
@@
The ranking loss used by perturbed optimization method described in Section 3.3 is convex and Lipschitz with respect to its parameters. Since the parameters correspond to the ratings themselves, the loss is also convex with respect to the parameters. Hence, assuming we restrict the ratings to a compact, convex set, there exists a global minimum that gradient descent via Fenchel-Young gradients is guaranteed to approach assuming standard learning rate conditions (e.g., square-summable, not summable).
4 Empirical Evaluation
We run experiments to demonstrate a number of properties of interest. In particular, we are interested in understanding how closely the rankings obtained from SCO approximate those obtained via Kemeny-Young, and aim to provide a comparison to outcomes returned by perturbed optimizers and Elo using several sources of data described below.
Example data.
@@ 4 Empirical Evaluation
PrefLib data.
These are examples from the voting literature on Wikipedia and on real data from elections, sports analytics, and others from PrefLib (Mattei and Walsh, 2013). Note that we restrict ourselves to the strictly-ordered incomplete (SOI) and strictly-ordered complete (SOC) data types in PrefLib, yielding a total of 12,680 preference profiles.
Synthetic evaluation data.
The synthetic data is generated to resemble those coming from online matching data, such as from a gaming site or tournament, based on TrueSkill (Herbrich et al., 2006). Agents’ true skill values are normally distributed and contests (matches) between them are generated such that each agent’s individual performance is stochastic with mean centered at their skill level. The outcome of each match-up is then a sorted list of each player’s performance in the match-up, equal to their true skill plus normally-distributed noise. Generated data allows us to mimic the structured sparsity present in real online game-play data but also to compare results to actual ground truth rankings.
Diplomacy game data.
This data set is our largest challenge problem: an anonymized agent-vs-agent data set of 7-player Diplomacy games played on the webDiplomacy web site (webdiplomacy.net) between 2008 and 2019 (Lanctot et al., 2023) with $m=52,958$ agents (players) and $n=31,049$ votes (games). Each game outcome is a strict order between seven players; the goal is to find a ranking over agents that minimizes the average Kendall-tau distance to all the votes. With these values, only $0.0011\%$ of the margin matrix entries are nonzero.
@@
For Elo ratings, the majorization-minorization algorithm of Hunter (used by bayeselo (Coulom, 2005)) is used to efficiently compute the best fit to the evaluation data (Hunter, 2004), and the SigmoidalProgramming package (Udell, 2020) to solve sigmoidal programs. Unless otherwise noted, by default we use gradient descent (Algorithm 1) to compute the SCO ratings, but also compare Fenchel-Young gradients and sigmoidal programming. Unless otherwise noted, $\theta_{min}=0$ and $\theta_{max}=100$. Full details such as specific hyperparameter values and additional experiments, please see Appendix D.
4.1 Warmup: Top-Ranking Condorcet Winners
Recall the example from Table 1. Since the overall pairwise win rates for $A$ and $C$ are equal (both beat every other agent in head-to-head match-up 60% of the time) (Lanctot et al., 2023), their gradients are identical and hence their Elo ratings are updated to the same value every iteration and hence converge to the equal values.
In contrast, SCO is designed to find the optimal ranking according to Definition 3, which top-ranks Condorcet winners when they exist. Consider the following 5-vote preference profile:

### v3

@@ 2.4. Learning-to-Rank
Another related field is that of learning-to-rank Liu (2011). The canonical example is that a user enters a keyword (query) and the problem is to retrieve the most relevant documents in a database, ranked by relevance to the keyword. There are several algorithms that learn to rank; Google’s PageRank, which powers their search engine, is one example. Our setting can be characterized by a learning-to-rank problem with a constant keyword (or no keyword) as there is no query. An important class of statistical models in this setting is random utility models Xia (2019). In random utility models, assessments are perceived as some ground truth assessment plus some noise.
@@
Perturbed optimizers Berthet et al. (2020); Blondel et al. (2020b) transform non-differentiable functions (such as sorting and ranking) into smoothed versions by adding noise; these perturbed functions can then be optimized by gradient descent. This is part of a growing effort to allow end-to-end training through discrete operators, using classical stochastic smoothing and perturbation approaches Gumbel (1954); Hazan et al. (2016). Included are optimal transport, clustering, dynamic time-warping and other dynamic programs (Cuturi, 2013; Cuturi and Blondel, 2017; Mensch and Blondel, 2018; Vlastelica et al., 2019; Paulus et al., 2020; Sander et al., 2023; Stewart et al., 2022) applied in fields such as computer vision, audio processing, biology, and physical simulators (Cordonnier et al., 2021; Kumar et al., 2021; Carr et al., 2021; Le Lidec et al., 2021; Baid et al., 2023; Llinares-López et al., 2023) and other optimization algorithms (Dubois-Taine et al., 2022).
3. Soft Condorcet Optimization
@@
An SCO rating is a numerical value representing an agent’s level of skill. SCO is closely related to several prior works: Elo Elo (1978), probabilistic ranking Diaconis (1988); Marden (1995); Alvo and Yu (2014), and perturbed optimizers Blondel et al. (2020b); Berthet et al. (2020). We elaborate on these relationships in Appendix D.
3.1. SCO Ratings and the Sigmoid Loss
@@
Let $\Pi(A)$ denote the set of permutations over alternatives $A$. We will call the data we work with “votes” to emphasize that we are working with (partial) rankings over alternatives, and borrow terminology from preferences and social choice. In particular, we view the entire dataset to be a collection of votes and so will refer to it as preference profile, []delimited-[]succeeds[$\succ$][ ≻ ]. Let $v\in[\succ]$ refer to each vote in the profile, where $v$ is a permutation over subsets of $A$, with length $|v|$. For a vote $v$, denote $v[i]$ as the alternative in position $i$ such that vote $v$ is represented as: $v[0]\succ v[1]\succ\cdots\succ v[|v|-1]$.
@@
Given some a profile []delimited-[]succeeds[$\succ$][ ≻ ] (i.e., set of votes), an optimal ranking minimizes the sum of the Kendall-tau distances
$R=argmin_{R\in\Pi(A)}\sum_{v\in[\succ]}K_{d}(v,R).$
@@
$L([\succ],A,V,\bm{\theta})=\sum_{v\in[\succ]}\sum_{(i,j)\in I_{2}(v)}D_{v}(\theta_{v[i]},\theta_{v[j]}),$ (3)
@@
Let []delimited-[]succeeds[$\succ$][ ≻ ] be the preference profile depicted in Table 1. We now show that the main loss function (Equation (3)) leads to same value as in Figure 1. The outer sum of Equation 3 enumerates the votes, which we will assume is in the same order as listed in Figure 1. The inner sum computes the Kendall-tau distance from the vote $v$ to the ranking $R$ induced by $\bm{\theta}$ (red exes in Figure 1). For the first vote $v=A\succ B\succ C$, the discrepancy function $D$ outputs 1 two times: once with pair $(i,j)=(0,2)$ and once with pair $(i,j)=(1,2)$ because the preferences between agents $(A,C)$ and agents $(B,C)$ disagree between $\bm{\theta}$ and $v$, so the inner sum for the first vote is 2. Similarly for the other votes: they correspond precisely to the same values as in Figure 1. Hence, the loss function (Equation (3)) is simply computing the sum of Kendall-tau distances between the ranking induced by $\bm{\theta}$ and all the votes.
@@
$\tilde{L}([\succ],A,V,\bm{\theta})=\sum_{v\in[\succ]}\sum_{(i,j)\in I_{2}(v)}\tilde{D}_{v}(\theta_{v[i]},\theta_{v[j]}).$ (6)
@@
Note that while we focus on the sigmoid loss as a soft approximation to Kendall-tau distance in this paper, the same approach can be used for other ranking distances that can be approximated by differentiable functions, such as Spearman’s footrule distance Diaconis and Graham (1977).
3.2. Sigmoid Loss Minimization
@@
3.2.1. Gradient Descent
The most straight-forward way is to apply gradient descent Hazan (2015): update ratings by following the gradient of the sigmoid loss, as shown in Algorithm 1. After applying the gradient to the ratings on line 1, the ratings may escape the bounded constraint space so we project them back. This is a straight-forward application of standard gradient descent (Hazan, 2015, Chapter 2). A common variant is stochastic gradient descent (SGD) which estimates the gradient by sampling subsets, i.e., “batches”, of the data set, computing the gradient using the sampled batch only. The standard $\ell_{2}$ projection step, Proj, projects the ratings back to the hypercube by clipping any ratings that fall outside the valid range $[\theta_{min},\theta_{max}]$.
We compute the gradient for a subset of the votes (i.e., batch $B\subseteq[\succ]$), $\nabla_{\bm{\theta}}\tilde{L}(B,\bm{\theta})$, from the batch loss $\tilde{L}(B,\bm{\theta})$, which resembles equation (6) but summed only over the votes in $B$ using the continuous $\tilde{D}_{v}$ from equation (5):
3.2.2. Sigmoidal Programming
A different way to compute SCO ratings is via sigmoidal programming Udell and Boyd (2014): solve for the (soft) optimum directly, i.e., find $\bm{\theta}^{*}$ that minimizes $\tilde{L}$ defined in equation 6. Note the $\tilde{L}$ can be rewritten in terms of the number of pairwise interactions between agents, quantified in the $\mathbf{N}$ matrix:
$\tilde{L}([\succ],A,V,\bm{\theta})$ $=$ $\sum_{a,b\in A\times A}N(a,b)\sigma(\theta_{b}-\theta_{a})$ (8)
which is a sum of sigmoidal functions $\sigma$ defined as functions which are strictly convex on domain $\theta_{b}-\theta_{a}\leq z$ and then strictly concave on $\theta_{b}-\theta_{a}\geq z$ (i.e.$z=0$). With a variable per difference in pair of ratings and appropriate constraints on variables and their feasible regions, the resulting optimization problem is known as a sigmoidal program which can be solved using a branch-and-bound algorithm Udell and Boyd (2014). A detailed construction is presented in Appendix C111 All appendices are available in the technical report version of the paper Lanctot et al. (2024).
Input: A preference profile []delimited-[]succeeds[$\succ$][ ≻ ]
Input: An initial parameter vector $\bm{\theta}^{0}=\left(\frac{\theta_{max}-\theta_{min}}{2}\right)\mathbf{1}$, where $\mathbf{1}$ is a vector of ones of length $|A|$
@@ 3.2.2. Sigmoidal Programming
2       $B\leftarrow$ $K$ votes sampled uniformly from []delimited-[]succeeds[$\succ$][ ≻ ]
3       Define $\nabla_{\bm{\theta}}\tilde{L}(B,\bm{\theta})$ based on equation (7).
4       $\bm{\theta}^{t}=\textsc{Proj}(\bm{\theta}^{t-1}-\alpha^{t}\nabla_{\bm{\theta}}\tilde{L}(B))$
@@
3.3. Fenchel-Young Loss Optimization
Here, we give an overview of the implementation of Fenchel-Young loss optimization; for more detail on the precise formulation and relationship perturbed optimizers, please see Appendix B.
In practice, Fenchel-Young loss optimization follows Algorithm 1, with a different definition of the loss and hence gradient on line 1. For a single vote $v$, a stochastic version of the gradient $\hat{g}_{v}$ can be computed: let $\bm{\theta}_{v}$ be the ratings for agents compared in $v$. Let $\mathbf{X}$ be a vector of Gumbel-distributed random variables of size $|v|$. Then $\tilde{\bm{\theta}}_{v}=\bm{\theta}_{v}+\sigma X$ is the perturbed ratings and let $\textsc{ArgSort}(-\tilde{\bm{\theta}}_{v})$ be the the indices of the elements that would sort the values. The Fenchel-Young gradient is then:
$\hat{g}_{v}=\textsc{ArgSort}(-\tilde{\bm{\theta}}_{v})-(0,1,\cdots,|v|-1),$ (9)
for all agents in $v$, and 0 otherwise. Inuitively, if the perturbed ranking obtained by sorting $\tilde{\bm{\theta}}_{v}$ would yield the same order of agents as in $v$, then the gradient for this vote would be zero. Otherwise, it is nonzero and each element of the gradient corresponds to the difference in rank position between the vote and perturbed ranking. Similarly to standard gradient descent, these gradients can be accumulated over batches $|B|\geq 1$ as $\nabla_{\bm{\theta}}\tilde{L}^{FY}(B,\bm{\theta})=\sum_{v\in B}\hat{g}_{v}$.
3.4. Theoretical Properties
@@
$\tilde{L}([\succ],A,V,\bm{\theta})$ $=$ $\sum_{a,b\in A\times A}N(a,b)\sigma(\theta_{b}-\theta_{a})\,,$ (10)
@@
Let $c\in A$ be the Condorcet winner. The loss $\tilde{L}$ as expressed in equation 8 is expressable in terms of a constant $K$ (that does not depend on $\bm{\theta}$) and a sum of sigmoids multiplied by coefficients from $\mathbf{M}$, by symmetry of the logistic function: $\sigma(x)=1-\sigma(-x)$. The second term is decomposable into contributions from comparisons to agent $c$, which is monotonically decreasing in $\theta_{c}$, and a sum which does not depend on $\theta_{c}$. As a result, increasing $\theta_{c}$ always decreases the loss, hence the minimum must correspond to $\theta^{*}_{c}=\theta_{max}$. ∎
@@
The ranking loss used by Fenchel-Young optimization method described in Section 3.3 is convex and Lipschitz with respect to its parameters. Since the parameters correspond to the ratings themselves, the loss is also convex with respect to the parameters. Hence, assuming we restrict the ratings to a compact, convex set, there exists a global minimum that gradient descent via Fenchel-Young gradients is guaranteed to approach assuming standard learning rate conditions (e.g., square-summable, not summable).
4. Empirical Evaluation
We run experiments to demonstrate a number of properties of interest. In particular, we are interested in understanding how closely the rankings obtained from SCO approximate those obtained via Kemeny-Young, compare to outcomes returned by perturbed optimizers and Elo using several sources of data:
Example data.:
@@ 4. Empirical Evaluation
PrefLib data.:
These are examples from the voting literature on Wikipedia and on real data from elections, sports analytics, and others from PrefLib Mattei and Walsh (2013). Note that we restrict ourselves to the strictly-ordered incomplete (SOI) and strictly-ordered complete (SOC) data types in PrefLib, yielding a total of 12,680 preference profiles.
Synthetic evaluation data.:
The synthetic data is generated to resemble those coming from online matching data, such as from a gaming site or tournament, based on TrueSkill Herbrich et al. (2006). Agents’ true skill values are normally distributed and contests (matches) between them are generated such that each agent’s individual performance is stochastic with mean centered at their skill level. The outcome of each match-up is then a sorted list of each player’s performance in the match-up, equal to their true skill plus normally-distributed noise. Generated data allows us to mimic the structured sparsity present in real online game-play data but also to compare results to actual ground truth rankings.
Diplomacy game data.:
This data set is our largest challenge problem: an anonymized agent-vs-agent data set of 7-player Diplomacy games played on the webDiplomacy web site (webdiplomacy.net) between 2008 and 2019 Lanctot et al. (2023) with $m=52,958$ agents (players) and $n=31,049$ votes (games). Each game outcome is a strict order between seven players; the goal is to find a ranking over agents that minimizes the average Kendall-tau distance to all the votes. With these values, only $0.0011\%$ of the margin matrix entries are nonzero.
@@
For Elo ratings, the majorization-minorization algorithm of Hunter (used by bayeselo Coulom (2005)) is used to efficiently compute the best fit to the evaluation data Hunter (2004), and the SigmoidalProgramming package Udell (2020) to solve sigmoidal programs. By default we use gradient descent to minimize the sigmoid loss (Algorithm 1) to compute the SCO ratings, but also compare Fenchel-Young gradients and sigmoidal programming. Unless otherwise noted, $\theta_{min}=0$ and $\theta_{max}=100$. Full details such as specific hyperparameter values, please see Appendix E. We also show two additional experiments in Appendix E: one which shows SCO used to approximate a Bayesian posterior over rating and one that shows SCO’s online performance.
4.1. Warmup: Top-Ranking Condorcet Winners
Lanctot et al. showed that Elo assigns the same rating to agent $A$ and $C$ in the example in Table 1, despite $A$ not being a Condorcet winner Lanctot et al. (2023). In contrast, SCO is designed to find the optimal ranking according to Definition 3, which top-ranks Condorcet winners when they exist. Consider the following 5-vote preference profile: